# ORCID ID: ]https://orcid.org/$0$-0002-9154-1878

### v1

@@ III.2 Antiferromagnetic interactions
Figure 5: Specific heat capacities (first line), sublattice magnetizations (second line), and the energy distribution functions near phase transition on a $24\times 24$ lattice (third line) for antiferromagnetic interaction ($J=1$) and representative values of $U$ ($U=-0.6,-0.4,-0.3,0.3$, from the left) at $\mu=0$.
@@
The case of antiferromagnetic interaction $J=1$ (see Fig. 4) is more complex, with a phase transition occurring only from the long-range antiferromagnetic (AF) phase, which, for $\mu=0$, occurs from $U=-0.5$ to $U=0$ (see $k_{B}T$-$U$ phase diagram for $\mu=0$ in Fig. 4). This AF phase is not an antiferromagnet in an usual meaning. In the AF phase, we found an alternate arrangement of magnetic moments in two sublattices (e.g., $A$ and $B$ sublattices, which form the hexagonal lattice; $m_{A}=-m_{B}$), whereas in the third sublattice (e.g., sublattice $C$) an average magnetization is zero ($m_{C}=0$, the direction of magnetic moment in a single realization is random), see snapshots in Fig. 4 (the labels of sublattices can be interchanged cyclically, cf. [29, 31]). In particular, for the half-filling case ($\mu=0$), this AF phase consists of the $C$ sublattice of a combination of empty and double-occupied sites (with compensated magnetic moments) occurring with equal probability.
@@
On the phase diagram, we also identified the SO phase, which exhibits only short-range ordering. Compared to mean-field results [32, 31], the spin configurations in this phase have similar combinations of (,,)($\uparrow$,$\uparrow$,$\downarrow$)( ↑ , ↑ , ↓ ) and (,,)($\uparrow$,$\downarrow$,$\downarrow$)( ↑ , ↓ , ↓ ) blocks. In contrary to the AF-NO transition line, the absence of long-range ordering does not lead to occurrence of phase transitions and the crossover from the SO phase to the NO phase is found.
@@
Similarly to the previous case, the position of the phase transitions and the crossover lines has been determined based on specific heat capacity and sublattice magnetizations, which are shown for representative values of $U$ in Fig. 5 (for $\mu=0$). The first column shows the typical behavior of the NO phase (for $U=-0.6$), characterized by zero sublattice magnetizations and a specific heat capacity curve that is not scaled with $L$. For $U\approx-0.4$, we observed the typical $\lambda$ singularity in the specific heat capacity curve, and the energy distribution functions indicate a second-order phase transition. Increasing $U$ gives rise to another peak in specific heat capacity, which is represented on the plot for $U=-0.2$. One can see that only the low-temperature peak scales (i.e., goes to infinity) with increasing lattice size $L$. At this temperature ($k_{B}T\approx 0.06$), we also observe a rapid change in sublattice magnetizations, which reveals the presence of a phase transition of the second order. On the other hand, the high-temperature peak (at $k_{B}T\approx 0.10$ for $U=-0.2$) shows a broad maximum with a slight shift as $L$ increases. Between these two peaks we observe fluctuations of sublattice magnetization, which are caused by the presence of the short-range ordering (and antiferromagnetic interactions causing frustration in the system). Similar behavior at the crossover from the non-ordered (NO) to short-range order (SO) phase with decreasing temperature is found also for $U>0$ (presented for $U=0.3$).
@@
Let us now proceed to the discussion of effects of nonzero chemical potential (a case of finite particle doping) and its influence on the behavior of the system, which is illustrated for representative values of $U$ in the $k_{B}T-\mu$ phase diagrams shown in Fig. 4. It should be noted that the antiferromagnetic AF phase occurring for $\mu\neq 0$ preferred empty sites in the $C$ sublattice for $\mu<0$ ($n<1$) and double-occupied sites for $\mu>0$ ($n>1$). Similar to $\mu=0$, the phase transitions occur only if the AF phase is present in the ground state. An increase in the chemical potential $|\mu|$ leads to an almost constant transition temperature for $U\approx-0.2$, or, alternatively, to initial increase of the transition temperature near $\mu\approx 0$ for $U=-0.1$. It drops to zero at the edge of the AF phase for $U=0$. The maximal AF-NO transition temperature moves to $\mu\neq 0$ with increasing $U$. Conversely, increasing $|\mu|$ lowers monotonously the temperature of the crossover between the NO phase and SO phase for all values of $U$. Note also that, for $\mu\neq 0$ and $U=1$ the AF phase occurs in finite temperatures even it does not exist for $\mu=0$ (such cases occur for $U>0$). It is also worth to mention that the maximal temperature of the AF phase occurrence is almost not dependent on $\mu$ and $U$ and it equals $k_{B}T\approx 0.06$, approximately.
@@
Figure 6: Specific heat capacities (first line), sublattice magnetizations (second line), number of particles (third row), and the energy distribution functions near phase transition on a $24\times 24$ lattice (fourth line) for antiferromagnetic interaction ($J=1$) and representative values of $U$ and $\mu$ (from the left: $U=-0.2$ and $\mu=-0.20,-0.15,-0.05$); $U=1.0$ and $\mu=-0.80,-0.60,-0.40$.
@@
Analogously to the previous case, in Fig. 6, we illustrate several physical quantities as a function of the temperature for representative values of $U$ ($U=-0.2$ and $U=1.0$) and $\mu$, which are used to construct the phase diagrams discussed above. The first and fourth columns (for $U=-0.2$ and $\mu=-0.2$ or $U=1.0$ and $\mu=-0.8$) represent the dependencies in the NO phase, where no phase transition occurs. For $k_{B}T\rightarrow 0$ the concentration $n\rightarrow 0$ in the NO phase with an agreement with exact ground state results [31]. The second column (for $U=-0.2$ and $\mu=-0.15$) shows the direct phase transition from the AF to the NO phase, characterized by a single peak in the specific heat capacity curve, a rapid change in sublattice magnetizations, and a number of particles approaching $n=2/3$ (and concentrations in the sublattices $n_{A}=n_{B}=1$, $n_{C}=0$) at the ground state. The third and fifth columns (for $U=-0.2$ and $\mu=-0.05$ or $U=1.0$ and $\mu=-0.60$) show specific heat capacity curves with two peaks: the low-temperature peak corresponds to the phase transition from the AF phase to the SO phase, while the high-temperature peak corresponds to the crossover between the SO state and the NO phases. It can be observed that the energy distribution function near the AF-SO transition temperatures exhibits a single peak, indicating that the phase transitions are of second order. The last column (for $U=1.0$ and $\mu=-0.40$) illustrates the case with the short-ordered ground state, where $n\rightarrow 1$ if $k_{B}T\rightarrow 0$ and sublattice magnetizations fluctuate strongly up to the crossover to the NO phase.
@@ IV Conclusion
In this work we investigated the localized fermion system on a triangular lattice described by the extended Hubbard model with onsite Coulomb $U$ and nearest-neighbor Ising-like magnetic $J$ interactions by means of the Monte carlo simulations in the grand canonical ensemble. By examining various quantities, such as specific heat capacity, sublattice magnetizations, number of particles and energy distribution functions, we construct comprehensive phase diagrams across varying temperatures and values of $U$ and chemical potential ($\mu$).
@@
The results for antiferromagnetic case ($J>0$) are very distinct from that obtained for $J<0$ case. This is because, in this case, the model exhibits geometrical frustration and thus the findings are strongly affected by the frustration. The region of the occurrence of the ordered AF phase is strongly reduced with only second-order transition from the AF phase (which is not an usual antiferromagnet). For $\mu=0$, the AF phase exist only for $U$ in the range of $-0.5<U<0$. For $U>0$ the AF phase can occur only for $\mu\neq 0$. The maximal temperature for the AF phase occurrence is $k_{B}T\approx 0.06$ and it is almost independent on $U$ or $\mu$.
@@
Finally, one should note that the considered system described by hamiltonian (1) is not as far from the reality as one can expect. In particular, the arrangement of magnetic moments in the triangular lattice has been observed in various materials, e.g., in Li2MnTeO6 [54], KCeO2 [55], KCeS2 [56], Ba2MnTeO6 [57, 58], Ce3Cu and Pr3Cu [59], Co1/3TaS2 [60], as well as in CsCeSe2 [61]. Moreover, also ultra-cold atomic gases give the opportunity of realization of systems with different geometries [62, 63, 64]. Because these systems have controllable parameters, they are perfect to study fundamental phenomena and allow to simulate different models in various regimes of parameters [65, 66, 24, 27, 67, 68]. The geometry of the lattice can be experimentally changed by different spatial arrangement of laser beams. In particular, it is possible to create the triangular lattice [69, 70, 71]. These facts create the future opportunity of the experimental validation of the predictions presented in the current work.

### v2

@@ III.2 Antiferromagnetic interactions
Figure 6: Specific heat capacities (first line), sublattice magnetizations (second line), and the energy distribution functions near phase transition on a $24\times 24$ lattice (third line) for antiferromagnetic interaction ($J=1$) and representative values of $U$ (U=0.6,0.4,0.3,0.3U=-0.6,-0.4,-$0.3,$ from the left) at $\mu=0$.
@@
The case of antiferromagnetic interaction $J=1$ (see Fig. 5) is more complex, with a phase transition occurring only from the long-range antiferromagnetic (AF) phase, which, for $\mu=0$, occurs from U=0.5U=-0.5 to $U=0$ (see $k_{B}T$-$U$ phase diagram for $\mu=0$ in Fig. 5). This AF phase is not an antiferromagnet in an usual meaning. In the AF phase, we found an alternate arrangement of magnetic moments in two sublattices (e.g., $A$ and $B$ sublattices, which form the hexagonal lattice; $m_{A}=-m_{B}$), whereas in the third sublattice (e.g., sublattice $C$) an average magnetization is zero ($m_{C}=0$, the direction of magnetic moment in a single realization is random), see snapshots in Fig. 5 (the labels of sublattices can be interchanged cyclically, cf. [29, 31]). In particular, for the half-filling case ($\mu=0$), this AF phase consists of the $C$ sublattice being a combination of empty and double-occupied sites (with compensated magnetic moments) occurring with equal probability.
@@
On the phase diagram, we also identified the SO state, which exhibits only short-range ordering. Compared to mean-field results [32, 31], the spin configurations in this phase have similar combinations of $(\uparrow,\uparrow,\downarrow)$ and $(\uparrow,\downarrow,\downarrow)$ blocks. In contrary to the AF-NO transition line, the absence of long-range ordering does not lead to occurrence of phase transitions and the crossover from the SO state to the NO phase is found. It is consistent with a fact that the SO state is not a new phase and it is just a state inside the NO phase. However, we explicitly distinguish the region of the SO state occurrence on the phase diagrams for $J>0$.
@@
Similarly to the previous case, the position of the phase transitions and the crossover lines has been determined based on specific heat capacity and sublattice magnetizations, which are shown for representative values of $U$ in Fig. 6 (for $\mu=0$). The first column shows the typical behavior of the NO phase (for U=0.6U=-0.6), characterized by zero sublattice magnetizations and a specific heat capacity curve that is not scaled with $L$. For $U\approx-0.4$, we observed the typical $\lambda$ singularity in the specific heat capacity curve, and the energy distribution functions indicate a second-order phase transition. Increasing $U$ gives rise to another peak in specific heat capacity, which is represented on the plot for U=0.2U=-0.2. One can see that only the low-temperature peak scales (i.e., goes to infinity) with increasing lattice size $L$. At this temperature ($k_{B}T\approx 0.06$), we also observe a rapid change in sublattice magnetizations, which reveals the presence of a phase transition of the second order. On the other hand, the high-temperature peak (at $k_{B}T\approx 0.10$ for U=0.2U=-0.2) shows a broad maximum with a slight shift as $L$ increases. Between these two peaks we observe fluctuations of sublattice magnetization, which are caused by the presence of the short-range ordering (and antiferromagnetic interactions causing frustration in the system). Similar behavior at the crossover from the non-ordered (NO) to short-range order (SO) phase with decreasing temperature is found also for $U>0$ (presented for $U=0.3$).
@@
Let us now proceed to the discussion of effects of nonzero chemical potential (a case of finite particle doping) and its influence on the behavior of the system, which is illustrated for representative values of $U$ in the $k_{B}T-\mu$ phase diagrams shown in Fig. 5. It should be noted that the antiferromagnetic AF phase occurring for $\mu\neq 0$ preferred empty sites in the $C$ sublattice for $\mu<0$ ($n<1$) and double-occupied sites for $\mu>0$ ($n>1$). Similar to $\mu=0$, the phase transitions occur only if the AF phase is present in the ground state. An increase in the chemical potential $|\mu|$ leads to an almost constant transition temperature for $U\approx-0.2$, or, alternatively, to initial increase of the transition temperature near $\mu\approx 0$ for U=0.1U=-0.1. It drops to zero at the edge of the AF phase for $U=0$. The maximal AF-NO transition temperature moves to $\mu\neq 0$ with increasing $U$. Conversely, increasing $|\mu|$ lowers monotonously the temperature of the crossover between the NO phase and SO state for all values of $U$. Note also that, for $\mu\neq 0$ and $U=1$ the AF phase occurs in finite temperatures even it does not exist for $\mu=0$ (such cases occur for $U>0$). It is also worth to mention that the maximal temperature of the AF phase occurrence is almost not dependent on $\mu$ and $U$ and it equals $k_{B}T\approx 0.06$, approximately.
@@
Figure 7: Specific heat capacities (first line), sublattice magnetizations (second line), number of particles (third row), and the energy distribution functions near phase transition on a $24\times 24$ lattice (fourth line) for antiferromagnetic interaction ($J=1$) and representative values of $U$ and $\mu$ (from the left: U=0.2U=-0.2 and $\mu=-0.15,-0.05$; $U=1.0$ and $\mu=-0.60,-0.40$).
@@
Analogously to the previous case, in Fig. 7, we illustrate several physical quantities as a function of the temperature for representative values of $U$ (U=0.2U=-0.2 and $U=1.0$) and $\mu$, which are used to construct the phase diagrams discussed above. The first column (for U=0.2U=-0.2 and $\mu=-0.15$) shows the direct phase transition from the AF to the NO phase, characterized by a single peak in the specific heat capacity curve, a rapid change in sublattice magnetizations, and a number of particles approaching $n=2/3$ (and concentrations in the sublattices $n_{A}=n_{B}=1$, $n_{C}=0$) at the ground state. The second and third columns (for U=0.2U=-0.2 and $\mu=-0.05$ or $U=1.0$ and $\mu=-0.60$) show specific heat capacity curves with two peaks: the low-temperature peak corresponds to the phase transition from the AF phase to the SO state (inside the NO phase), while the high-temperature peak corresponds to the crossover between the SO state and the NO phases. It can be observed that the energy distribution function near the AF-SO transition temperatures exhibits a single peak, indicating that the phase transitions are of second order. The last column (for $U=1.0$ and $\mu=-0.40$) illustrates the case with the short-ordered ground state, where $n\rightarrow 1$ if $k_{B}T\rightarrow 0$ and sublattice magnetizations fluctuate strongly up to the crossover to the NO phase.
@@ IV Conclusion
In this work we investigated the localized fermion system on a triangular lattice described by the extended Hubbard model with onsite Coulomb $U$ and nearest-neighbor Ising-like magnetic $J$ interactions by means of the Monte carlo simulations in the grand canonical ensemble. The investigated model (1) is a non-trivial extension of the well-known Ising model. By examining various quantities, such as specific heat capacity, sublattice magnetizations, number of particles and energy distribution functions, we construct comprehensive phase diagrams for both signs of $J$ interaction across varying temperatures and values of $U$ and chemical potential ($\mu$).
@@
The results for antiferromagnetic case ($J>0$) are very distinct from that obtained for $J<0$ case. This is because, in this case, the model exhibits geometrical frustration and thus the findings are strongly affected by the frustration. The region of the occurrence of the ordered AF phase is strongly reduced with only second-order transition from the AF phase (which is not an usual antiferromagnet). For $\mu=0$, the AF phase exist only for $U$ in the range of 0.5<U<0-0.5<U<0. For $U>0$ the AF phase can occur only for $\mu\neq 0$. The maximal temperature for the AF phase occurrence is $k_{B}T\approx 0.06$ and it is almost independent on $U$ or $\mu$. Note also that two sublattices, from which single occupied sites form the antiferromagnetic pattern in the AF phase defined here (with $n\approx 2/3$ or $4/3$), create the honeycomb lattice. For the Ising model on the honeycomb lattice ($z=3$), the critical temperature for the ordered phase is found as $kT_{c}/(2|J|)=0.253$ [11].
@@
Finally, one should note that the considered system described by hamiltonian (1) is not as far from the reality as one can expect. In particular, the arrangement of magnetic moments in the triangular lattice has been observed in various materials, e.g., in Li2MnTeO6 [54], KCeO2 [55], KCeS2 [56], Ba2MnTeO6 [57, 58], Ce3Cu and Pr3Cu [59], Co1/3TaS2 [60], as well as in CsCeSe2 [61]. Moreover, also ultra-cold atomic gases give the opportunity of realization of systems with different geometries [62, 63, 64]. Because these systems have controllable parameters, they are perfect to study fundamental phenomena and allow to simulate different models in various regimes of parameters [65, 66, 24, 27, 67, 68]. The geometry of the lattice can be experimentally changed by different spatial arrangement of laser beams. In particular, it is possible to create the triangular lattice [69, 70, 71]. As neglecting of the single-particle hopping might be considered as an artificial one, it can be recognized as the first step for description of the magnetism in the mentioned systems. These facts create the future opportunity of the experimental validation of the predictions presented in the current work.